# Maximum Independent Set; Hardness Parameter; Neutral Atoms; Rydberg Blockade; QuEra’s Aquila; Analog Quantum Computing; Adiabatic Schedules; Counterdiabatic Driving

### v2

@@ 2.2 Counterdiabatic protocol
The limited coherence time of the quantum state in experiments is a major source of error in adiabatic computing. Well-established methods to overcome such limitations include the shortcuts to adiabadicity, among which is the counterdiabatic (CD) driving [36]. The idea of the CD driving is to add a gauge potential to the adiabatic Hamiltonian such that the transitions between eigenstates are suppressed during the evolution. However, the calculation of the exact gauge potential is a very difficult task in general, and these mathematical terms may not be always adapted to an experimental implementation. Nevertheless, by neglecting the interactions of the Rydberg Hamiltonian, one is able to find a suitable gauge potential which utilizes the phase of the Rabi drive [28]. In this study, we propose to generalize the adiabatic schedules of this so-called Analog Counterdiabatic Quantum Computing (ACQC) protocol as follows:
@@
where $\theta(t)$ is any monotonic smooth function that ranges from 00 to $\pi$; here we set $\theta_{\Omega,\Delta}(t)=\pi\,t/t_{\text{f}}$. Moreover, based on the variational procedure described in [27], we improve on the ACQC protocol by including the Rydberg interactions in the calculation of the CD terms, thus adapting the gauge potential to each different graph, see Appendix B. The detuning is not altered by this operation, whereas the Rabi drive becomes:
@@
and where $0\leq\nu\leq C_{6}/a^{6}$ is an additional free parameter that interpolates between the standard and the generalized gauge potentials. In the latter equation, the variables $T_{1}$ and $T_{2}$ depend on the interaction terms $V_{ij}$ and are therefore specific to the graph to be solved. As expected, the CD contribution vanishes for $t_{\text{f}}\to\infty$ since $\dot{\Omega}_{\text{ad}}$ and $\dot{\Delta}_{\text{ad}}$ tend to zero in this limit.
@@
Similarly as for the piecewise linear protocol, we optimize the schedule parameters $(\Delta_{\text{i}},\Delta_{\text{f}},\nu)$ for all graphs of the dataset, see Fig. 4. The interpolating coefficient $\nu$ tends to a strictly positive value for the hardest graphs, indicating that the generalized gauge potential plays an important role in the CD protocol. Note that the easy graphs exhibit optimized values of $\Delta_{\text{f}}$ that are smaller than the statistically determined constant $\Delta_{\text{LB}}$. This is due to the fact that the exact lower bounds on the detuning are much smaller in that case, see Fig. 11 in the appendix. The shape of the CD phase is highly nontrivial, and we observe a strong shift toward higher values of the detuning for hard MIS instances, see Fig. 5. Finally, we also tried to further improve the results by considering more evolved $\theta$ functions—in particular, to mimic the skewed shape of the Rabi amplitude in Fig. 2—but no significant gain was found in this case.
Hardware implementation
In Aquila, the Rabi amplitude and the detuning must be defined as piecewise linear and not smooth functions, while the phase must be piecewise constant. However, the step size can be as small as 0.05 µs, which results in 80 intervals for the longest coherent evolution. The schedules are therefore well approximated in the quantum hardware, but one could slightly improve the experimental results by taking into account this hardware capability in the numerical simulations.
@@
Figure 4: Optimum parameters for the CD schedules. The initial detuning (top) reaches the upper bound $-\delta_{\Delta}$ for several graphs but tends to a lower value $\Delta_{\text{i}}\approx-\Delta_{\text{UB}}/10$ for the hardest graphs. The final detuning (bottom) also increases with the hardness parameter and quickly reaches the limit $\Delta_{\text{UB}}$. The coefficient $\nu$ (not shown) displays a similar behavior and tends to $\nu=4.30$ in the current setting.
@@
Figure 5: Optimized counterdiabatic schedules with 3 free variables. Unlike the detuning (bottom), the Rabi amplitude (top) and phase (middle) depend not only on $\mathcal{HP}$ but also on $t_{\text{f}}$ and on the graph-dependent variables $T_{1}$ and $T_{2}$. Here, the curves are displayed for $t_{\text{f}}=1$ and for the mean values of the dataset $\overline{T}_{1}\approx 0.93$ and $\overline{T}_{2}\approx 1.46$. The light green area represents the variation due solely to $T_{1}$ and $T_{2}$ for $\mathcal{HP}=0.5$.
2.3 Results
@@
In Fig. 6, we compare the probability $P_{\text{MIS}}$ to reach the solution for 4 protocols: the piecewise linear and ACQC protocols defined in [28] which serve as benchmarks, and their generalized and hardness-dependent counterparts described in the previous sections. Note that $a=5.5$ for the benchmark protocols but $a=5$ in our study. The $\mathcal{HP}$-linear protocol (with 6 degrees of freedom) yields the best results for all graphs, and in particular for the hard ones with a tenfold improvement over the benchmark.
@@
Graph-to-graph comparisons are illustrated in Fig. 7. As for the optimized schedule parameters, we observe that the values of $P_{\text{MIS}}$ depend mostly on the hardness parameter and not on the size of the graphs. For instance, the success probabilities for the $\mathcal{HP}$-linear protocol range from 0.59 ($\mathcal{HP}=4.20$) to 0.99 ($\mathcal{HP}=0.46$) for $N=10$, and from 0.11 ($\mathcal{HP}=9.71$) to 0.98 ($\mathcal{HP}=0.52$) for $N=17$.
@@
Figure 6: Box plot of the probability to reach the MIS depending on the hardness of the 500 graphs of the dataset. The piecewise linear (blue) and ACQC (orange) protocols serve as benchmarks. For hard graphs, the generalized $\mathcal{HP}$-ACQC (green) and $\mathcal{HP}$-linear (red) protocols developed in this study lead to an average success probability that is improved by approximately 150% and 1 000%, respectively.
@@
Figure 7: Graph-to-graph comparisons of $P_{\text{MIS}}$. Left: the $\mathcal{HP}$-linear protocol developed in this study leads to a 3.6-fold mean improvement on the linear protocol described in [28]. Right: except for rare instances, the hardness-dependent piecewise linear protocol always beats its ACQC counterpart. Note that $P_{\text{MIS}}$ depends strongly on $\mathcal{HP}$ but not on $N$ (upward or downward triangle), with low sucess probabilities for the hard graphs (maroon) and nearly perfect results for the easy ones (green).
3 Experiments
@@
Based on thorough numerical simulations, we have shown in the previous section that hardness-dependent schedules lead to high probabilities to reach the MIS solution. Now, we extend these results to the hardware and to graphs that are not part of the dataset used in the optimization. First, we consider a small set of “toy graphs” that we can still simulate classically to ensure that the various protocols are correctly run in Aquila. Then, we apply the models to large graphs made up of 137 vertices to verify that the hardness parameter and not the size of the graph is indeed the most relevant property when solving the MIS problem.
Measurement shots
@@
For all examples, we use 720 shots (repetitions of the adiabatic evolution) to get enough data for a statistical comparison of the methods. We choose this specific value because it is a multiple of 9, 12, and 16, which are the maximum number of copies of the small graphs ($3\times 3$ and $3\times 4$ lattices with $a=5$ and $a=5.5$) that fit and can be run simultaneously in the hardware. We consider only the shots in which the lattice is correctly filled by atoms. The typical probability of failing to occupy a site in Aquila is $\epsilon_{\text{fill}}=0.007$ [21, Sec. 1.5], and in practice we get $684\pm 7$ and $380\pm 14$ valid shots for the small and the large graphs, respectively.
3.1 Toy graphs
@@
In Fig. 8, we display the probability to reach the MIS solution for the 11 toy graphs defined in [29]. As in that article, we set $t_{\text{f}}=0.7$ so that the probabilities of success are distributed around 50%. We observe a good agreement between numerical simulations and the corresponding results obtained on hardware, especially for the benchmark protocols. A possible reason for the optimized schedules to yield outcomes that are slightly lower than expected may be the choice of a smaller lattice spacing or the larger gradient of $\Omega$ at the end of the evolution; many other sources of error could explain these differences, see [21, Sec. 1.5] for an exhaustive list. Nevertheless, the $\mathcal{HP}$-linear protocol still performs better on the hardware than the linear and CD benchmarks, with a threefold and a twofold mean improvement, respectively.
@@
Figure 8: Probability to reach the MIS solution for the 11 toy graphs. The rectangle bars are numerical results and the error bars are experimental results (95% confidence intervals). The ranking of the 4 studied protocols is the same as in Fig. 6, with $\mathcal{HP}\in[2,3]$ here. The black horizontal line represents the average success probability obtained by the Bayesian optimization in [29, Fig. 9], which performs better than the benchmarks (blue and orange) but not as well as the hardness-dependent protocols (green and red).
3.2 Large graphs
@@
We now turn to large graphs, which are of much higher interest in real applications. We consider three of the four graphs studied in [29, Fig. 10], which consist of 137 vertices and whose hardness parameters extend over three orders of magnitude ($\mathcal{HP}\approx$ 14, 126, and 1435). In addition, we generate one very easy graph of the same size to span yet another order of magnitude ($\mathcal{HP}\approx$ 1.4), see Fig. 9. The hardness parameter of these large graphs is efficiently computed by advanced classical algorithms, namely by the GenericTensorNetworks Julia library [37, 38].
@@
Since the MIS solutions and consequently the hardness parameters are not known a priori, we apply our optimized linear schedules in the limit $\mathcal{HP}\to\infty$. Moreover, we set $t_{\text{f}}=t_{\text{max}}=4$ so that the probability to find a MIS solution is as high as possible. Finally, we post-process all experimental data such that the measured outcomes become valid maximal independent sets. To this end, we use the greedy procedure described in [19, Sec. S2.3]: first, we remove any of two neighboring vertices until the sets become independent. Second, we include a vertex at any free site with no neighbor until the sets are maximal. The same procedure applied to an empty configuration is used as a classical benchmark, which highlights the relative gain obtained with the quantum protocols [21, Sec. 6].
@@
The results are summarized in Fig. 9: the MIS solution is never found by the classical greedy algorithm and is obtained by the standard linear protocol only for the easiest graph. On the contrary, the $\mathcal{HP}$-linear protocol leads to the solution for all but the hardest graph, which indicates a clear improvement over the existing protocols.
4 Discussion and outlook
@@
First, a statistical investigation of the conditions for which the final quantum state encodes the MIS solution has revealed that the final detuning should lie in a narrow interval. To our knowledge, these are the tightest bounds ever mentioned. We believe that choosing a final detuning in that interval would improve the results of many other studies, either in terms of success probability or number of optimization steps. In fact, applying adiabatic protocols in which the final ground state does not correspond to the sought solution can only lead to poorly performing algorithms.
@@
Second, the individual optimization of numerous small graphs has shown that the optimum parameters of the schedules depend mostly on the hardness parameter and only marginally on the size of the graph. Through extensive classical simulations of small systems, we have generated adiabatic schedules that perform extremely well also when applied to much larger graphs; only the hardest instances are not solved by our protocol. Obviously, improvements in the hardware would greatly help in this regard, especially if it permits a longer coherent evolution time. Since the quantum technology evolves rapidly, this could be the case soon. However, because of the combinatorial nature of the hardness parameter, very hard MIS problems exist only in sufficiently large graphs, so that our numerical study should be extended to larger graphs to address this issue. We believe that improved schedules could be easily found by considering slightly larger systems, for instance up to about 25 vertices, or even larger ones using tensor network techniques for the simulations whenever appropriate [39].
@@
We have focused our work on two families of schedules, namely piecewise linear and smooth counterdiabatic functions. The phase of the Rabi drive is nonconstant only in the latter case, and we have developed a generalized version of the ACQC protocol [28] to make it specific to each graph. However, when correctly parametrized, the somehow simpler $\mathcal{HP}$-linear protocol performs better in almost all numerical simulations. This difference is even more pronounced in experiments, as the smooth CD schedules have to be converted into piecewise linear or constant functions. We have used a maximum number of 6 free parameters in the optimization, but more general functions may be investigated since the method described in this work applies to any kind of schedules. Moreover, it would be interesting to develop models of optimized schedules that do not depend only on the hardness parameter but also on other global properties of the graphs, e.g., the mean degree or the mean eccentricity of the vertices. Based on machine learning, such additional predictors could in fact improve the fitting of the schedule parameters.
@@
We think that hybrid classical-quantum algorithms could take advantage of our protocols and, more generally, of precomputed optimized schedules: instead of searching in the huge space of all parameters, one could consider lower-dimensional meaningful ansatzes, such as the $\mathcal{HP}$-parametrization of the schedules or the principal component of the residuals of the fitted parameters.
@@
Finally, we believe that the piecewise linear protocols developed in this study should serve as a new benchmark for any future adiabatic algorithm that aims to solve the MIS problem with Rydberg atoms. In particular, the most basic version of this protocol with only 4 free parameters and in the limit of infinite hardness admits a very simple mathematical description, see Fig. 2, and it leads to relatively good results regardless of the specific graph to be solved.
@@
The supporting data for this article are openly available on GitHub [40].
[…]
@@
Figure 9: Experiments on large unit disk graphs. Top: four graphs with 137 vertices each and whose hardness parameters span four orders of magnitude; the three hardest graphs are taken from [29]. In shades of blue, the probability $\braket{\hat{n}_{i}}$ to measure the $i$-th atom in the Rydberg state |$\ket{1}$ at the end of the $\mathcal{HP}$-linear protocol. Due to the high degeneracy of the MIS, some vertices are equally measured in the ground or excited state. Bottom: experimental results (after a minimal post-processing) to find an independent set of a given size, with the MIS solution at the rightmost bin. The error bars represent the 95% confidence intervals. The classical greedy algorithm (gray) fails to find the solution in all cases. The benchmark linear protocol (blue) reaches the MIS in the easiest case but fails for any harder instance. Finally, the $\mathcal{HP}$-linear protocol is able to find the MIS except for the hardest graph; the corresponding values of $P_{\text{MIS}}$ are indicated in the subtitles. The success probabilities for the two CD protocols (not shown) lie between those of the two linear protocols.
[…]
@@ Appendix A Ground state encoding of the maximum independent set problem
Given a simple undirected graph $\mathcal{G}=(V,E)$ with vertex set $V$ and edges set $E$, a subset $S\subseteq V$ is called independent if no pair of vertices in $S$ are connected by an edge in $E$. In the maximum independent set (MIS) problem, we want to find the independent set of largest cardinality. Denoting by $x_{i}\in\{0,1\}$ the absence or the presence of the vertex $i=1,\ldots,N$ in $S$, it is proven that the MIS solution minimizes the cost function [41, 42]
@@
if $0<\Delta<V_{ij}$ for all $ij\in E$. In analog quantum computing with arrays of Rydberg atoms, $\mathcal{G}$ is a unit disk graph that connects the nearest and next-nearest neighbors of a square lattice. Neglecting the interactions between farther atoms and for a non-planar graph (i.e., when crossing diagonals belong to it), the MIS condition becomes:

### v3

@@ II.2 Counterdiabatic protocol
The limited coherence time of the quantum state in the hardware is a major source of error in adiabatic computing. Well-established methods to overcome such limitations include the shortcuts to adiabaticity, among which is the CD driving [44]. The idea of the CD driving is to add a gauge potential to the adiabatic Hamiltonian such that the transitions between eigenstates are suppressed during the evolution. However, the calculation of the exact gauge potential is very difficult in general, and these mathematical terms may not be always adapted to an implementation in the quantum hardware. Nevertheless, by ignoring the interactions of the Rydberg Hamiltonian, one is able to find a suitable gauge potential that utilizes the phase of the Rabi drive [32]. In this study, we propose to generalize the adiabatic schedules of this so-called analog counterdiabatic quantum computing (ACQC) protocol as follows:
@@
where $\theta(t)$ is any monotonic smooth function that ranges from 00 to $\pi$; here we set $\theta_{\Omega,\Delta}(t)=\pi t/t_{\text{f}}$. Moreover, on the basis of the variational procedure described in Ref. [31], we improve on the ACQC protocol by including the Rydberg interactions in the calculation of the CD terms, thus adapting the gauge potential to each different graph; see Appendix B. The detuning is not altered by this operation, whereas the Rabi drive becomes
@@
and where $0\leq\nu\leq V_{0}$ is an additional free parameter that interpolates between the standard gauge potential and the generalized gauge potential; see Appendix B.2.2. In the latter equation, the variables $T_{1}$ and $T_{2}$ depend on the interaction terms $V_{ij}$ and are therefore specific to the graph to be solved. As expected, the CD contribution vanishes for $t_{\text{f}}\to\infty$ since both $\dot{\Omega}_{\text{ad}}$ and $\dot{\Delta}_{\text{ad}}$ tend to zero in this limit.
@@
Similarly as for the piecewise linear protocol, we optimize the schedule parameters $(\Delta_{\text{i}},\Delta_{\text{f}},\nu)$ for all graphs of the dataset; see Fig. 5. The interpolating coefficient $\nu$ tends to a strictly positive value for the hardest graphs, indicating that the generalized gauge potential plays an important role in the CD protocol. Note that the easy graphs exhibit optimized values of $\Delta_{\text{f}}$ that are smaller than the statistically determined constant $\Delta_{\text{LB}}$. This is because the exact lower bounds on the detuning are much smaller in that case; see Fig. 12 in Appendix A.2. The shape of the CD phase is highly nontrivial, and we observe a strong shift toward higher values of the detuning for hard MIS instances; see Fig. 6. Finally, we also tried to further improve the results by considering more evolved $\theta$ functions—in particular, to mimic the skewed shape of the Rabi amplitude in Fig. 2—but no significant gain was found in this case.
@@
Figure 5: Optimum parameters for the CD schedules. The initial detuning (top) reaches the upper bound $-\delta_{\Delta}$ for several graphs but tends to a lower value $\Delta_{\text{i}}\approx-\Delta_{\text{UB}}/10$ for the hardest graphs. The final detuning (bottom) also increases with the hardness parameter and quickly reaches the limit $\Delta_{\text{UB}}$. The coefficient $\nu$ (not shown) displays a similar behavior and tends to approximately $V_{0}/80$ in the current setting.
@@
Figure 6: Optimized counterdiabatic schedules with three free variables. Unlike the detuning (bottom), the Rabi amplitude (top) and phase (middle) depend not only on $\mathbb{H}$ but also on $t_{\text{f}}$ and on the graph-dependent variables $T_{1}$ and $T_{2}$. Here the curves are displayed for $t_{\text{f}}=1$ and for the mean values of the dataset $\overline{T}_{1}\approx 0.93$ and $\overline{T}_{2}\approx 1.46$. The light-green area represents the variation solely due to $T_{1}$ and $T_{2}$ for $\mathbb{H}=0.5$.
II.3 Results
@@
In Fig. 7, we compare the probability $P_{\text{MIS}}$ to reach the solution for four protocols: the piecewise linear and ACQC protocols defined in Ref. [32], which serve as benchmarks, and their generalized and hardness-dependent counterparts described in the previous sections. Note that $a=5.5$ for the benchmark protocols but $a=5$ in our study. The $\mathbb{H}$-linear protocol (with six degrees of freedom) yields the best results for all graphs, and in particular for the hard ones with a tenfold improvement over the benchmark.
@@
Graph-to-graph comparisons are illustrated in Fig. 8. As for the optimized schedule parameters, we observe that the values of $P_{\text{MIS}}$ depend mostly on the hardness parameter and not on the size of the graphs. For instance, the success probabilities for the $\mathbb{H}$-linear protocol range from 0.59 ($\mathbb{H}=4.20$) to 0.99 ($\mathbb{H}=0.46$) for $N=10$, and from 0.11 ($\mathbb{H}=9.71$) to 0.98 ($\mathbb{H}=0.52$) for $N=17$.
@@
Figure 7: Box plot of the probability of reaching the MIS depending on the hardness of the 500 graphs of the dataset. The piecewise linear (blue) and ACQC (orange) protocols serve as benchmarks. For hard graphs, the generalized $\mathbb{H}$-ACQC (green) and $\mathbb{H}$-linear (red) protocols developed in this study lead to an average success probability that is increased by approximately 150% and 1 000%, respectively.
@@
Figure 8: Graph-to-graph comparisons of $P_{\text{MIS}}$. Left: The $\mathbb{H}$-linear protocol developed in this study leads to a 3.6-fold mean improvement on the linear protocol described in Ref. [32]. Right: Except for rare instances, the hardness-dependent piecewise linear protocol always beats its ACQC counterpart. Note that $P_{\text{MIS}}$ depends strongly on $\mathbb{H}$ but not on $N$ (upward or downward triangles), with low success probabilities for the hard graphs (maroon) and nearly perfect results for the easy ones (green).
III Cloud quantum computing demonstrations
@@
On the basis of thorough numerical simulations, we showed in the previous section that hardness-dependent schedules lead to high probabilities of reaching the MIS solution. We first extend these results to the hardware and to graphs that are not part of the dataset used in the optimization. First, we consider a small set of “toy graphs” that we can still simulate classically to ensure that the various protocols are correctly run in Aquila. Then we apply the models to large graphs made up of 137 vertices to verify that the hardness parameter and not the size of the graph is indeed the most relevant property when solving the MIS problem.
@@
For all examples, we use 720 shots (repetitions of the adiabatic evolution) to get enough data for a statistical comparison of the methods. We choose this specific value because it is a multiple of 9, 12, and 16, which are the maximum number of copies of the small graphs ($3\times 3$ and $3\times 4$ lattices with $a=5$ and $a=5.5$) that fit and can be run simultaneously in the hardware. We consider only the shots in which the lattice is correctly filled by atoms. The typical probability of failing to occupy a site in Aquila is $0.007$ (see Sec. 1.5 in Ref. [25]), and in practice we get $684\pm 7$ and $380\pm 14$ valid shots for the small and the large graphs, respectively.
III.1 Toy graphs
@@
In Fig. 9, we display the probability of reaching the MIS solution for the 11 toy graphs defined in [33]. As in that article, we set $t_{\text{f}}=0.7$ so that the probabilities of success are distributed around 50%. We observe good agreement between numerical simulations and the corresponding results obtained on hardware, especially for the benchmark protocols. A possible reason for the optimized schedules yielding outcomes that are slightly lower than expected may be the choice of a smaller lattice spacing or the larger gradient of $\Omega$ at the end of the evolution; many other sources of error could explain these differences—see Sec. 1.5 in [25] for an exhaustive list. Nevertheless, the $\mathbb{H}$-linear protocol still performs better on the hardware than the linear and CD benchmarks, with a threefold mean improvement and a twofold mean improvement, respectively.
@@
Figure 9: Probability of reaching the MIS solution for the 11 toy graphs. The rectangular bars are numerical results and the error bars are the results of the demonstration (95% confidence intervals). The ranking of the four protocols studied is the same as in Fig. 7, with $\mathbb{H}\in[2,3]$ here. The horizontal black line represents the average success probability obtained by the Bayesian optimization in Ref. Finzgar2024 (see Fig. 9 therein), which performs better than the benchmarks (blue and orange) but not as well as the hardness-dependent protocols (green and red).
III.2 Large graphs
@@
We now turn to large graphs, which are of much greater interest in real applications. We consider three of the four graphs studied in Ref. [33] (see Fig. 10 therein), which consist of 137 vertices and whose hardness parameters extend over 3 orders of magnitude ($\mathbb{H}\approx$ 14, 126, and 1435). In addition, we generate one very easy graph of the same size to span yet another order of magnitude ($\mathbb{H}\approx$ 1.4); see Fig. 10. In these examples, the hardness parameter can still be computed by an advanced classical algorithm; namely, by the GenericTensorNetworks Julia library [45, 46]. However, in general, this parameter cannot be considered to be known beforehand as it depends on the MIS solution, which is the very target of the demonstrations. Therefore, we apply the optimized schedules in the limit $\mathbb{H}\to\infty$ in what follows, and we focus on the piecewise linear protocols as they are expected to yield the best results.
@@
Contrarily to the small graphs, we now use the full coherent time of the hardware, i.e. $t_{\text{f}}=t_{\text{max}}=4$. We want the probability of finding a MIS solution to be as high as possible for the hardest instances. Moreover, we postprocess all data of the demonstrations such that the measured outcomes become valid maximal independent sets. To this end, we use the greedy procedure described in Sec.  S2.3 in Ref. [19]: First, we remove any of two neighboring vertices until the sets become independent. Second, we include a vertex at any free site with no neighbor until the sets are maximal. The same procedure applied to an empty configuration is used as a classical benchmark, which highlights the relative gain obtained with the quantum protocols [25, 36].
@@
The results are summarized in Fig. 10: The MIS solution is never found by the classical greedy algorithm and is obtained by the standard linear protocol only for the easiest graph. In contrast, the $\mathbb{H}$-linear protocol leads to the solution for all but the hardest graph, which indicates a clear improvement over the existing protocols.
IV Discussion and outlook
@@
First, a statistical investigation of the conditions for which the final quantum state encodes the MIS solution has revealed that the final detuning should lie in a narrow interval. To our knowledge, these are the tightest bounds ever mentioned. We believe that choosing a final detuning in that interval would improve the results of many other studies, either in terms of success probability or in terms of the number of optimization steps. Indeed, the application of adiabatic protocols in which the final ground state does not correspond to the sought solution can only lead to poorly performing algorithms.
@@
Second, the individual optimization of numerous small graphs has shown that the optimum parameters of the schedules depend mostly on the hardness parameter and only marginally on the size of the graph. Through extensive classical simulations of small systems, we have generated adiabatic schedules that perform extremely well also when applied to much larger graphs; only the hardest instances are not solved by our protocol. Obviously, improvements in the hardware would greatly help in this regard, especially if they lead to a longer coherent evolution time. Since quantum technology is evolving rapidly, this could be the case soon. However, because of the combinatorial nature of the hardness parameter, very hard MIS problems exist only in sufficiently large graphs, so our numerical study should be extended to larger graphs to address this issue. We believe that improved schedules could be easily found by consideration of slightly larger systems, with up to about 25 vertices, or even larger ones with the use of tensor network techniques for the simulations whenever appropriate [47].
@@
We have focused our work on two families of schedules; namely, piecewise linear and smooth counterdiabatic functions. The phase of the Rabi drive is nonconstant only in the latter case, and we have developed a generalized version of the ACQC protocol [32] to make it specific to each graph. However, when correctly parametrized, the somehow simpler $\mathbb{H}$-linear protocol performs better in almost all numerical simulations. This difference is even more pronounced in demonstrations, as the smooth CD schedules have to be converted into piecewise linear or constant functions. We have used a maximum number of six free parameters in the optimization, but more general functions may be investigated since the method described in this work applies to any kind of schedule. Moreover, it would be interesting to develop models of optimized schedules that do not depend only on the hardness parameter but that depend also on other global properties of the graphs, e.g., the mean degree or the mean eccentricity of the vertices. On the basis of machine learning, such additional predictors could improve the fitting of the schedule parameters.
@@
We think that hybrid classical-quantum algorithms could take advantage of our protocols and, more generally, of precomputed optimized schedules: instead of searching in the huge space of all parameters, one could consider lower-dimensional meaningful ansatzes, such as the $\mathbb{H}$ parametrization of the schedules or the principal component of the residuals of the fitted parameters.
@@
Finally, we believe that the piecewise linear protocols developed in this study should serve as a benchmark for any future adiabatic algorithm that aims to solve the MIS problem with Rydberg atoms. In particular, the most basic version of this protocol with only four free parameters and in the limit of infinite hardness admits a very simple mathematical description—see Fig. 2—and it leads to relatively good results regardless of the specific graph to be solved.
Data availability
@@
The data that support the findings of this article are openly available [48].
@@
Figure 10: Demonstrations for large unit disk graphs. Top: Four graphs with 137 vertices each and whose hardness parameters span 4 orders of magnitude; the three hardest graphs are taken from [33]. In shades of blue, the probability $\braket{\hat{n}_{i}}$ to measure the $i$th atom in the Rydberg state |$\ket{1}$ at the end of the $\mathbb{H}$-linear protocol. Because of the high degeneracy of the MIS, some vertices are equally measured in the ground state or the excited state. Bottom: Results of the demonstrations (after a minimal postprocessing) to find an independent set of a given size, with the MIS solution at the rightmost bin. The error bars represent the 95% confidence intervals. The classical greedy algorithm (gray) fails to find the solution in all cases. The benchmark linear protocol (blue) reaches the MIS in the easiest case but fails for any harder instance. Finally, the $\mathbb{H}$-linear protocol is able to find the MIS except for the hardest graph; the corresponding values of $P_{\text{MIS}}$ are indicated. The success probabilities for the two CD protocols (not shown) lie between those of the two linear protocols.
@@
Given a simple undirected graph $\mathcal{G}=(V,E)$ with vertex set $V$ and edge set $E$, a subset $S\subseteq V$ is called independent if no pair of vertices in $S$ are connected by an edge in $E$. In the MIS problem, we want to find the independent set of largest cardinality. If we denote by $x_{i}\in\{0,1\}$ the absence or the presence of the vertex $i=1,\ldots,N$ in $S$, it is proven that the MIS solution minimizes the cost function [49, 50]
@@
if $0<\Delta<V_{ij}$ for all $ij\in E$. In analog quantum computing with arrays of Rydberg atoms, $\mathcal{G}$ is a unit disk graph that connects the nearest and next-nearest neighbors of a square lattice. If we ignore the interactions between farther atoms, for a non-planar graph (i.e., when crossing diagonals belong to it), the MIS condition becomes